lected. (Assume 1 Cu atom per formula unit.) mass of the test tube and unknown sample mass of test tube without the unknown sample mass of filter paper mass of filter paper and copper (Assume 1 Cu atohfaunknown copper compound from this data that a student col 14.603 g 13.437 g 0.633 g 1.096 g
Answer #1

o Given. Copper atem per formula uvut = →Nero = 1-1668-0. 46st man X, =0.703 mous 아X = 14-6039 . 13.437 molar naus 아 07038 fo